Macguffing have a demand function of Qdx=70-p and a supply function of Qsx=2p+10 determine the quantity supplied is 0?


1
Expert's answer
2021-03-08T07:24:47-0500

"Q_{dx}=Q_{sx}""70-P=2P+10"

"70-10=3P"

"60=3P"

"P=60\/3=20"

"Q_{sx}=2\times20+10=50"